Show that IP(m + k,m) = P(m + nm + 1) (m + 1)
Answer #1

T o ( n +1) & P ( m+ K, m) = P(mth, m+1) since P (mrk, m) = (mrk)! mi to Comp 22 + (rmy 22 + Coup) --- Afghan Onti)! +(mi2]!Toge Na (m+2)! +m+22! tomt3)? -- (min-131 (na) = (m+2)! (opt 3s ) + Cmta?! + ---Company?? fm 1.3)? [CM J + --- mampul - mt4)